# A note on the triple product property for finite groups with abelian normal subgroups of prime index

### v2

@@ Abstract.
Three non-empty subsets $S,T,U$ of a group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, $t,t^{\prime}\in T$, $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. In this case $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ is called the size of the triple. If $G$ is a finite group the triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [4, Conjecture 7.6]. This note proves a version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any finite group which contains an abelian normal subgroup of prime index $p$.
@@
fast matrix multiplication, triple product property, finite groups, TPP triple, TPP capacity, subgroup TPP capacity, TPP ratio, subgroup TPP ratio
@@
The author is not affiliated with any institution. The work contained in this note was originally completed in 2013-2014, and for this the author is grateful for the assistance of Prof. Sarah Hart, Birkbeck, University of London, and Dr. Peter M. Neumann, The Queen’s College, Oxford.
@@ 1. Introduction
TPP triples were first introduced by H. Cohn and C. Umans in 2003 to study the complexity of fast matrix multiplication [3]. Further applications and developments of this approach were described in [2, 4, 5]. However this note pursues a combinatorial problem for TPP triples in finite groups introduced by Peter M. Neumann in [6]. This involves two quantities. The first quantity, called the subgroup TPP ratio of $G$, is defined as
@@
It is natural to seek best upper bounds for $\beta(G)$ and $\beta_{0}(G)$ (equivalently, for $\rho(G)$ and $\rho_{0}(G)$). It was shown by Cohn and Umans that if $G$ is a dihedral group then $\rho(G)\geq\frac{4}{3}$ [3]. P. Neumann derived a general upper bound for $\beta(G)$ [6, Corollary 3.2] that
@@
using the following observation [6, Observation 3.1]:
@@
In 2012 Hedtke and Murthy conjectured [4, Table 1 and Conjectures 7.5-7.6], based on tables of TPP triple data in groups with cyclic subgroups of index $2$ of order up to $32$, obtained via exhaustive search algorithms in these groups implemented using the GAP computer algebra system, that
@@
This note focuses on subgroup TPP ratio $\rho_{0}$ (equivalently, subgroup TPP capacity $\beta_{0}$), and proves a generalisation of the conjecture for $\rho_{0}$ for all groups with abelian normal subgroups of prime index. First, a few relevant properties and characterisations of TPP triples are stated, followed by proofs of two technical results needed to establish the main result.
@@ 2. Elementary Properties of TPP Triples
[6, p. 234] A TPP triple $(S,T,U)$ of $G$ is called basic if
@@
For non-empty subsets $X,Y\subseteq G$ the quotient set $Q(X,Y)$ of $X$ and $Y$ is the set $Q(X,Y)=XY^{-1}=\{xy^{-1}|x\in X,y\in Y\}$. If $X=Y$ then $Q(X,Y)=Q(X)=XX^{-1}$.
@@
[4, Theorem 3.1] For non-empty subsets $S,T,U\subseteq G$ the triple $(S,T,U)$ is a TPP triple if and only if
@@
    If $\pi\in\textrm{Sym}\{S,T,U\}$, i.e. $\pi$ is a permutation of $\{S,T,U\}$, then $(S^{\pi},T^{\pi},U^{\pi})$ is a TPP triple of $G$ (permutation invariance) [3, Lemma 2.1].
@@
    If $a,b,c,d\in G$ are any elements then $(dSa,dTb,dUc)$ is a TPP triple of $G$ (translation invariance) [6, Observation 2.1].
@@
All future references to TPP triples will be to basic TPP triples, unless otherwise stated.
The following two results are useful to note.
@@
Given a TPP triple $(S,T,U)$ of a group $G$ and a subgroup $H\leq G$ there is a subgroup TPP triple $(S\cap H,T\cap H,U\cap H)$ of $H$.
@@ 3. Technical Results
(2) Let $S\leq G$. If $H\trianglelefteq G$ then using the Second Isomorphism Theorem $SH\leq G$ and $\overline{S}=SH/H\leq G/H$, where $SH/H\cong S/(S\cap H)$
@@
Let $G$ be a group, $H\trianglelefteq G$ an abelian subgroup, $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$, and $(S,T,U)$ a subgroup TPP triple of $G$. Let $\overline{S},\overline{T},\overline{U}\leq G/H$ be the $H$-supports of $S,T,U$ respectively, and $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, $U_{0}:=U\cap H$.
(1) If $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$ then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ are cosets of $S_{0}T_{0}U_{0}$ in $H$, where $S_{x}:=S\cap xH$, $T_{x}:=T\cap xH$, $U_{y}:=U\cap yH$. Additionally these cosets are non-trivial if $xH,yH\neq H$ and distinct if $xH\neq yH$.
(2) If $U_{0}=U\cap H\trianglelefteq G$ and $xH,yH\in\overline{S}\cap\overline{T}$ such that $xH\neq yH$ then $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, where $S_{x}:=S\cap xH$, $T_{x}:=T\cap xH$.
@@
Let $H$ be as given. We may assume $H<G$, otherwise there is nothing to prove. Note that for any $gH=Hg\in G/H$ its inverse is $(gH)^{-1}=g^{-1}H=Hg^{-1}$, and if $X_{g}\subseteq gH$ is any non-empty subset its inverse is $X_{g}^{-1}\subseteq g^{-1}H=Hg^{-1}$. Recall that $(S_{0},T_{0},U_{0})$ is a subgroup TPP triple of $H$, and, as $H$ is abelian, $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. Also recall that, as subgroups with the TPP, $S,T,U$ satisfy $S\cap TU=T\cap U=\{1\}$.
@@
(1) Let $S_{x},T_{x}$ be defined as above. As $S_{x},T_{x}\neq\emptyset$ elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$ can be chosen. Then $s_{x}^{-1}t_{x}\in H$ and $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}\subseteq H$. This shows that $S_{x}^{-1}T_{x}U_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and $|S_{x}^{-1}T_{x}U_{0}|=|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. Clearly if $xH\neq H$, that is, if $x\in G\backslash H$, then $S_{x}^{-1}T_{x}U_{0}$ is a non-trivial coset of $S_{0}T_{0}U_{0}$ in $H$. Now, also assume that $yH\in\overline{S}\cap\overline{U}$ and let $S_{y}:=S\cap yH$ and $U_{y}:=U\cap yH$. By transposing $T$ and $U$, and $x$ and $y$, in the argument above, it can be shown that $S_{y}^{-1}U_{y}T_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and, furthermore, a non-trivial coset if $yH\neq H$.
@@
Assume $xH\neq yH$. Suppose $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}\neq\emptyset$. Then there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $u_{y}\in U_{y}$, $t_{0}\in T_{0}$, and $u_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}u_{y}t_{0}$. We can rewrite this as $s_{x}s_{y}^{-1}u_{y}t_{0}u_{0}^{-1}t_{x}^{-1}=1$. Since $t_{0},u_{0}$ commute (as elements of the abelian group $H$), this can be rewritten as $s_{x}s_{y}^{-1}u_{y}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. The TPP implies that $u_{y}=u_{0}$ and $t_{x}=t_{0}$, a contradiction. So $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$, which proves part (1).
@@
(2) Assume $U_{0}=U\cap H\trianglelefteq G$ and $xH,yH\in\overline{S}\cap\overline{T}$ such that $xH\neq yH$. From above $S_{x}^{-1}T_{x}U_{0},S_{y}^{-1}T_{y}U_{0}\subseteq H$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$. Suppose there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $t_{y}\in T_{y}$, $u_{0},\widetilde{u}_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}t_{y}\widetilde{u}_{0}$. We can rewrite this as $st_{y}\bar{u}_{0}t_{x}^{-1}=1$, where $s=s_{x}s_{y}^{-1}$ and $\bar{u}_{0}=\widetilde{u}_{0}u_{0}^{-1}$. By the normality of $U_{0}$ in $G$ there is a $\hat{u}_{0}\in U_{0}$ such that $\bar{u}_{0}t_{x}^{-1}=t_{x}^{-1}\hat{u}_{0}$. So $st_{y}\bar{u}_{0}t_{x}^{-1}=st_{y}t_{x}^{-1}\hat{u}_{0}=1$. The TPP implies that $t_{x}=t_{y}$, a contradiction. Thus $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, which proves part (2). ∎

### v5

@@ Abstract.
Three non-empty subsets $S,T,U$ of a group $G$ are said to satisfy the triple product property (TPP) if, for elements $s,s^{\prime}\in S$, and $t,t^{\prime}\in T$, and $u,u^{\prime}\in U$, the equation $s^{\prime}s^{-1}t^{\prime}t^{-1}u^{\prime}u^{-1}=1$ holds if and only if $s=s^{\prime}$, $t=t^{\prime}$, $u=u^{\prime}$. If this is the case then $(S,T,U)$ is called a TPP triple of $G$ and $|S||T||U|$ the size of the triple. If $G$ is a finite group the triple product ratio of $G$ can be defined as the quantity $\rho(G):=\frac{\beta(G)}{|G|}$, where $\beta(G)$ is the largest size of a TPP triple of $G$, and a special case of this, the subgroup triple product ratio, is the quantity $\rho_{0}(G):=\frac{\beta_{0}(G)}{|G|}$, where $\beta_{0}(G)$ is the largest size of a TPP triple of $G$ composed only of subgroups. There is a conjecture that $\rho(G)\leq\frac{4}{3}$ if $G$ contains a cyclic subgroup of index $2$ [5, Conjecture 7.6]. This note proves a more general version of this conjecture for subgroups by showing that $\rho_{0}(G)\leq\frac{p^{2}}{2p-1}$ if $G$ is any finite group which contains an abelian normal subgroup of prime index $p$, and discusses its implications for $\rho$ for groups with cyclic normal subgroups of prime index, based on the known data for $\rho$ in such groups of small order.
@@
triple product property, finite groups, TPP triple, TPP capacity, subgroup TPP capacity, TPP ratio, subgroup TPP ratio, fast matrix multiplication
@@
The author is unaffiliated. With the exception of some minor updates, all the work contained in this note was originally completed in 2013-2014, and for this the author is grateful for the assistance of Prof. Sarah Hart, Birkbeck, University of London, and Dr. Peter M. Neumann, The Queen’s College, Oxford.
@@ 1. Introduction
TPP triples were first introduced by H. Cohn and C. Umans in 2003 to study the complexity of fast matrix multiplication [3]. Further applications and developments of this approach were described in [2, 5, 6]. However this note pursues the combinatorial problems for TPP triples in finite groups introduced by Peter M. Neumann in [7]. These involve two quantities. The first quantity, called the subgroup TPP ratio of $G$, is defined as
@@
It is natural to seek best upper bounds for $\beta(G)$ and $\beta_{0}(G)$ (equivalently, for $\rho(G)$ and $\rho_{0}(G)$). It was shown by Cohn and Umans that if $G$ is a dihedral group then $\rho(G)\geq\frac{4}{3}$ [3]. P. Neumann derived a general upper bound for $\beta(G)$ [7, Corollary 3.2] that
@@
using the following observation [7, Observation 3.1]:
@@
In 2012 Hedtke and Murthy conjectured [5, Table 1 and Conjectures 7.5-7.6], based on tables of TPP triple data in groups with cyclic subgroups of index $2$ of order up to $32$, obtained via exhaustive search algorithms in these groups implemented using the GAP computer algebra system, that
@@
This note focuses on subgroup TPP ratio $\rho_{0}$ (equivalently, subgroup TPP capacity $\beta_{0}$), and proves a generalisation of the conjecture for $\rho_{0}$ for all groups with abelian normal subgroups of prime index. First, a few relevant properties and characterisations of TPP triples are stated, followed by proofs of two technical results needed to establish the main result. Following the proof of the main result some implications for $\rho$ for groups with cyclic normal subgroups of prime index are discussed, based on known data for $\rho$ in such groups.
@@ 2. Elementary Properties of TPP Triples
[7, p. 234] A TPP triple $(S,T,U)$ of $G$ is called basic if
@@
For non-empty subsets $X,Y\subseteq G$ the quotient set $Q(X,Y)$ of $X$ and $Y$ is the subset $Q(X,Y)=XY^{-1}=\{xy^{-1}|x\in X,y\in Y\}\subseteq G$. If $X=Y$ then $Q(X,Y)=Q(X)=XX^{-1}$.
@@
[5, Theorem 3.1] Three non-empty subsets $S,T,U\subseteq G$ satisfy the TPP if and only if
@@
    If $\pi\in S_{3}$, i.e. $\pi$ is a permutation of $\{S,T,U\}$, then $(S^{\pi},T^{\pi},U^{\pi})$ is a TPP triple of $G$ (permutation invariance) [3, Lemma 2.1].
@@
    If $a,b,c,d\in G$ are any elements then $(dSa,dTb,dUc)$ is a TPP triple of $G$ (translation invariance) [7, Observation 2.1].
@@
All further references to TPP triples will be to basic TPP triples, unless otherwise stated.
The following two elementary properties are useful to note, and will be used at several points.
@@
Given a TPP triple $(S,T,U)$ of a group $G$ and a subgroup $H\leq G$ there is a TPP triple $(S\cap H,T\cap H,U\cap H)$ of $H$.
@@
Part (2) of Proposition 2.5 means that only non-abelian groups can realise TPP triples of non-trivial size.
@@ 3. Technical Results
(2) Let $S\leq G$. If $H\trianglelefteq G$ then using the Second Isomorphism Theorem $SH\leq G$ and $\overline{S}=SH/H\leq G/H$, where $SH/H\cong S/(S\cap H)$.
@@
Let $G$ be a group, $H\trianglelefteq G$ an abelian subgroup, $G/H=\{gH\mid g\in G\}$ the quotient group of $G$ by $H$, and $(S,T,U)$ a subgroup TPP triple of $G$. Furthermore, let the $H$-supports $\overline{S}$, $\overline{T}$, $\overline{U}$ be defined as in Observation 3.2, where these are subgroups of $G/H$.
(1) If $xH,yH$ are any non-trivial cosets of $H$ such that $S_{x},S_{y}\in\overline{S}$, $T_{x}\in\overline{T}$ and $U_{y}\in\overline{U}$, and letting $S_{0}:=S\cap H$, $T_{0}:=T\cap H$, and $U_{0}:=U\cap H$, then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}U_{y}T_{0}$ are non-trivial cosets of $S_{0}T_{0}U_{0}$ in $H$, that is, $S_{x}^{-1}T_{x}U_{0},S_{y}^{-1}U_{y}T_{0}\neq S_{0}T_{0}U_{0}$, and also distinct, that is, $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$.
(2) If $xH,yH$ are any non-trivial and distinct cosets of $H$ then $S_{x}^{-1}T_{x}U_{0}$ and $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$.
@@
Let $H$ be as given ($H$ is abelian and normal in $G$), and $H<G$ can be assumed, otherwise there is nothing to prove. Several facts used in the proof are noted for convenience:
  • The inverse of any coset $xH=Hx\in G/H$ is $(xH)^{-1}=x^{-1}H=Hx^{-1}$, and the inverse of any non-empty subset $B_{x}\subseteq xH$ is $B_{x}^{-1}\subseteq x^{-1}H=Hx^{-1}$ (by the normality of $H$ in $G$).
  • If $x_{1},x_{2}\in xH$ are any coset elements then $x_{1}x_{2}^{-1}\in H$ (if $x_{1}=xh_{1},x_{2}=xh_{2}$ for some $h_{1},h_{2}\in H$ then $x_{1}x_{2}^{-1}=xh_{1}h_{2}^{-1}x^{-1}=xx^{-1}h_{3}=h_{3}\in H$ for some $h_{3}\in H$ (by the normality of $H$ in $G$).
  • Every permutation $(S^{\pi},T^{\pi},U^{\pi})$ of the TPP triple $(S,T,U)$, for $\pi\in S_{3}$, is also a TPP triple (permutation invariance, Observation 2.3).
  • The subgroups $S_{0},T_{0},U_{0}$ as defined above, which satisfy the TPP, also have the property that $S_{0}\trianglelefteq S$, $T_{0}\trianglelefteq T$, $U_{0}\trianglelefteq U$ and $SH/H\cong S/S_{0}$, $TH/H\cong T/T_{0}$, $UH/H\cong U/U_{0}$ (Second Isomorphism Theorem). This means that translates of the form $S_{0}s_{x}=s_{x}S_{0},T_{0}t_{y}=t_{y}T_{0},U_{0}u_{z}=u_{z}U_{0}$, for $s_{x}\in S\cap xH$, $t_{y}\in T\cap yH$, $u_{z}\in U\cap zH$ and $xH,yH,zH\in G/H$, are cosets of $S_{0},T_{0},U_{0}$ respectively in $S,T,U$ respectively which intersect with the cosets $xH,yH,zH$ respectively.
  • As $H$ is abelian, the subgroups $S_{0}$, $T_{0}$, and $U_{0}$, are also normal in $H$ and thus form a set product $S_{0}T_{0}U_{0}$ which is a subgroup of $H$ of order $|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$ (using the injective triple product map on $S_{0}\times T_{0}\times U_{0}$ into $H$ from Proposition 2.5).
@@
(1) Let $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$, and $S_{x},S_{y},T_{x},U_{y}$ be defined as above. Note that the sets $S_{x}^{-1}=(S\cap xH)^{-1}=S^{-1}\cap(xH)^{-1}=S\cap x^{-1}H$ and $S_{y}^{-1}=(S\cap yH)^{-1}=S\cap y^{-1}H$ and $S_{x}^{-1}T_{x},S_{y}^{-1}T_{y},S_{y}^{-1}U_{y}\subseteq H$ will also appear by calculation. As $S_{x},T_{x}\neq\emptyset$ elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$ can be chosen. Then, as $s_{x}^{-1}t_{x}\in H$ it follows that $s_{x}^{-1}t_{x}(S_{0}T_{0}U_{0})=S_{0}s_{x}^{-1}t_{x}T_{0}U_{0}=S_{x}^{-1}T_{x}U_{0}\subseteq H$, which shows that $S_{x}^{-1}T_{x}U_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$ of size $|S_{x}^{-1}T_{x}U_{0}|=|S_{0}T_{0}U_{0}|=|S_{0}||T_{0}||U_{0}|\leq|H|$. Additionally, if $xH\neq H$ (if $x\in G\backslash H$) then this is coset is non-trivial, i.e. $S_{x}^{-1}T_{x}U_{0}\neq S_{0}T_{0}U_{0}$. To see this, suppose there are elements $s_{x}\in S_{x}$, $t_{x}\in T_{x}$, $u_{0}\in U_{0}$ and $s_{0}\in S_{0}$, $t_{0}\in T_{0}$, $\widetilde{u}_{0}\in U_{0}$, such that $s_{x}^{-1}t_{x}u_{0}=s_{0}^{-1}t_{0}\widetilde{u}_{0}$. This can be rearranged as $s_{x}s_{0}^{-1}t_{0}\widetilde{u}_{0}u_{0}^{-1}t_{x}^{-1}=1$. As $t_{0}$ and $\widetilde{u}_{0}u_{0}^{-1}$ commute in $H$ this can be further rearranged as $s_{x}s_{0}^{-1}\widetilde{u}_{0}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $(S,U,T)$ implies that $s_{0}=s_{x}$ and $t_{0}=t_{x}$, a contradiction. By transposing $T$ and $U$, and $x$ and $y$, in the argument above, it can also be shown that $S_{y}^{-1}U_{y}T_{0}$ is a coset of $S_{0}T_{0}U_{0}$ in $H$, and, furthermore, a non-trivial coset if $yH\neq H$.
@@
Suppose $xH,yH\neq H$ and $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}\neq\emptyset$. Then there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $u_{y}\in U_{y}$, $t_{0}\in T_{0}$, and $u_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}u_{y}t_{0}$, or $s_{x}s_{y}^{-1}u_{y}t_{0}u_{0}^{-1}t_{x}^{-1}=1$. Since $t_{0}$ and $u_{0}$ commute in $H$ this can be rewritten as $s_{x}s_{y}^{-1}u_{y}u_{0}^{-1}t_{0}t_{x}^{-1}=1$. However the TPP for $S,U,T$ implies that $u_{y}=u_{0}$ and $t_{x}=t_{0}$, a contradiction. So $S_{x}^{-1}T_{x}U_{0}\cap S_{y}^{-1}U_{y}T_{0}=\emptyset$ if $xH\in\overline{S}\cap\overline{T}$ and $yH\in\overline{S}\cap\overline{U}$ such that $xH,yH\neq H$, which proves part (1).
@@
(2) Assume $U_{0}=U\cap H\trianglelefteq G$ and let $xH,yH\in\overline{S}\cap\overline{T}$ be such that $xh,yH\neq H$ and $xH\neq yH$. Suppose there are elements $s_{x}\in S_{x}$, $s_{y}\in S_{y}$, $t_{x}\in T_{x}$, $t_{y}\in T_{y}$, $u_{0},\widetilde{u}_{0}\in U_{0}$ such that $s_{x}^{-1}t_{x}u_{0}=s_{y}^{-1}t_{y}\widetilde{u}_{0}$, or, $st_{y}\bar{u}_{0}t_{x}^{-1}=1$, where $s=s_{x}s_{y}^{-1}$ and $\bar{u}_{0}=\widetilde{u}_{0}u_{0}^{-1}$. By the normality of $U_{0}$ in $G$ there is a $\hat{u}_{0}\in U_{0}$ such that $\bar{u}_{0}t_{x}^{-1}=t_{x}^{-1}\hat{u}_{0}$, which implies that $st_{y}\bar{u}_{0}t_{x}^{-1}=st_{y}t_{x}^{-1}\hat{u}_{0}=1$. However, the TPP for $S,T,U$ implies that $t_{x}=t_{y}$, a contradiction. Thus $S_{x}^{-1}T_{x}U_{0}$, $S_{y}^{-1}T_{y}U_{0}$ are distinct cosets of $S_{0}T_{0}U_{0}$ in $H$, which proves part (2). ∎
4. The Main Result